Circular No. 7000

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998dq IN NGC 6754 F. Patat and P. Leisy, European Southern Observatory (ESO); and

A. Piemonte, Pontificia Universidad Catolica, write: “We obtained a CCD spectrogram (10-min exposure, range 340–745 nm, resolution 2.6 nm) of SN 1998dq (cf. IAUC 6998) on Aug. 27.00 UT with the ESO 3.6-m telescope (+ EFOSC2) at La Silla. A full reduction of the spectrum shows a strong Si II 635.5-nm feature, a multitude of Fe II lines, and the presence of Ca H and K, allowing us to conclude that this a type-Ia supernova around maximum light and very similar to SN 1994D (Patat et al. 1996, MNRAS 278, 111). The expansion velocities, deduced from the absorption minima of Si II 635.5-nm and Ca H and K, are 12 000 and 17 000 km/s respectively. An unresolved absorption of Na I D is clearly visible at the parent galaxy’s rest-frame velocity (3257 km/s; Mathewson and Ford 1996, Ap.J. Suppl. 107, 97), as in the case of nearby SN 1998X. The measured total equivalent width of the Na I D doublet is ∼ 0.11 nm.”

XTE J2123–058 C. Zurita and J. Casares, Instituto de Astrofísica de Canarias (IAC),

report: “CCD R-band images of the x-ray transient XTE J2123–058 (IAUC 6955) have been obtained on Aug. 26 and 27 using the 1-m Optical Ground Station telescope owned by the European Space Agency and operated by the IAC at the Observatorio del Teide. The object has declined to an average magnitude of R = 21.50 ± 0.06 and the previously reported 1.4-mag orbital modulation has completely disappeared (IAUC 6993). Instead, our data suggest the dominance of the ellipsoidal modulation from the companion star: a double-humped variability with a peak-to-peak amplitude of 0.20 ± 0.16 mag. The current magnitude is consistent with pre-outburst estimates (IAUC 6957, 6993), indicating that the transient has probably reached quiescence. This represents the best opportunity so far to detect the companion star spectroscopically.”

COMET 9P/TEMPEL 1 P. Lamy, Laboratoire d’Astronomie Spatiale, and his team report the

detection of the nucleus of comet 9P with the Hubble Space Telescope (+ WFPC2) on 1997 Dec. 31 at Δ = 3.53 AU and r = 4.48 AU: “The slightly incomplete lightcurve indicates an elongated body having semi-axes a = 3.9 and b = 2.8 km (assuming a geometric albedo of 0.04) and a rotational period of ∼ 25 hr. No coma was detected.”

1998 August 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green